OpenStudy (daenio):

12. An undercover news reporter wanted to know what exactly Fast Food hamburgers were made of. It was found that 95% of the burgers were made from beef, the rest was made from a Mystery Meat! a) What is the probability that at least 18 out of 20 of hamburgers selected randomly are made of beef? b) How many hamburgers would you expect to be made from Mystery Meat, out of a sample of 120 hamburgers? I find this to be pretty tricky. Could u give me an idea on what I'm supposed to do, so I can figure out the rest?

OpenStudy (daenio):

@jim_thompson5910

OpenStudy (anonymous):

at least 18 out of 20 means 18 or 19 or 20 in reverse order \[P(X=20)=.95^{20}\] \[P(X=19)=20\times .95^{19}\times .05\] \[P(X=18)=\dbinom{20}{2}.95^{18}\times .05^2\]

OpenStudy (anonymous):

\[\dbinom{20}{2}=\frac{20\times 19}{2}=190\]

OpenStudy (anonymous):

add them up for your answer expected value is easier multiply \[120\times .95\]

OpenStudy (anonymous):

no that is wrong. expected value of mystery meat is \[.05\times 120\]

OpenStudy (daenio):

Thank you! :D

OpenStudy (daenio):

I'm confused. What's the answer to a)?

OpenStudy (daenio):

@jim_thompson5910